# Exact Learning Dynamics of In-Context Learning in Linear Transformers and Its Application to Non-Linear Transformers

### v1

@@ 4 Applications to Analysis of Non-linear Transformer Models
Probing Timescale Separation via Curvature. The solved loss dynamics (Eq. (14)) predict non-uniform learning rates across different spectral modes $\alpha$, leading to characteristic plateaus followed by cliffs in the loss curve as modes $\alpha$ converge according to their timescales $\tau_{\alpha}\propto 1/s_{\alpha}^{2}$. To empirically detect these points of rapid change in potentially complex loss curves, we examine the curvature of the loss autocorrelation function. High curvature points should correspond to transitions in the learning process. Calculating the second derivative of the loss autocorrelation indeed reveals prominent peaks corresponding to expected mode activations (Figure 4A). Complementarily, we analyze the curvature of the Frobenius norm of the active parameter matrices ($\|p_{2}(t)\|_{F}$ or $\|q_{1}(t)\|_{F}$). Since the parameters $p_{\alpha},q_{\alpha}$ evolve according to Eq. (12) with mode-dependent rates, their overall norm should also exhibit non-uniform changes. Peaks in the parameter norm curvature align well with those in the loss autocorrelation (Figure 4B), reinforcing that macroscopic observables reflect the underlying timescale separation predicted by the theory.
@@
Tracking Sequential Learning via Effective Rank. The exact solution $a_{\alpha}(t)$ (Eq. (13)) describes the sequential "activation" or learning of different spectral modes. To track this process empirically in the weight matrices of non-linear models, we utilize the effective rank [26], which measures the dimensionality or concentration of singular values ($s_{i}$) of a matrix $M$:
@@
A lower effective rank implies energy concentration in fewer singular modes. Motivated by the theoretical prediction that dominant modes (large $s_{\alpha}$) learn first (small $\tau_{\alpha}$), we expect the effective rank to initially decrease as the model focuses on these modes, and subsequently increase as subdominant modes are learned. To better visualize this sequential process, we introduce "marginalized effective rank". This involves computing the effective rank repeatedly, each time excluding the singular value component corresponding to the mode assumed to have converged earliest. Dips in these marginalized curves should indicate the time points where specific spectral components undergo significant learning (Figure 4C). As predicted by the sequential learning in Eq. (13), the effective rank initially dips, then recovers, and the marginalized curves reveal staggered learning across modes.
@@
We applied these analyses to attention-only non-linear transformers (1-4 layers), similar to those exhibiting ICL capabilities [18]. Figure 6 displays the marginalized effective rank for Output-Value (OV) and Query-Key (QK) matrices. Consistent with the timescale separation hypothesis, models capable of ICL (2+ layers) show characteristic dips in the effective rank curves, particularly for QK matrices, around the time ICL performance emerges (highlighted region). The full set of marginalized curves reveals multiple minima occurring sequentially, supporting the idea derived from our linear model that learning proceeds in stages along different spectral dimensions. In contrast, the 1-layer model, which fails at this ICL task, shows a monotonic increase in OV effective rank, suggesting undifferentiated learning or noise fitting rather than structured, sequential learning. These empirical results suggest that the sequential, mode-dependent learning predicted by our solvable model qualitatively extends to these non-linear settings and correlates with the emergence of functional capabilities like ICL.
Stability of Parameter Subspace. Our theoretical analysis separates the dynamics of parameter magnitudes (related to $a_{\alpha}=p_{\alpha}q_{\alpha}$) from the underlying parameter structure. The decoupled dynamics (Eq. (12)) allow for the possibility that the directions spanned by the singular vectors of weight matrices stabilize faster than the singular values reach their fixed points. To test this, we introduce a "Subspace Distance" metric, measuring the alignment between the subspace spanned by the singular vectors of a weight matrix $M(t)$ at time $t$ and its final state $M(\infty)$:
[…]
@@
This hypothesis is further tested by examining models exhibiting “grokking” in modular arithmetic tasks [25], where generalization is significantly delayed relative to training performance saturation. Figure 3B shows the characteristic delayed drop in test error. Analyzing the OV matrix dynamics, we observe that subspace stabilization, measured by the Subspace Distance (Figure 7C), is markedly delayed. It occurs significantly after the training loss has minimized and coincides closely with the sharp improvement in generalization (test error drop). The effective rank (Figure 7B) also shows dynamics linked to this delayed phase. This finding strongly suggests that grokking, in this context, is linked to a delayed convergence of the parameter subspace, providing a potential mechanistic explanation. Once the correct subspace is identified (albeit late), learning of magnitudes proceeds rapidly, leading to generalization.
@@ 5 Discussion
The learned computation at the fixed point (Eq. (11)) effectively implements the target linear map $W^{\mu}$ by applying an operator $\hat{\Sigma}_{x}[s^{\infty}(\Sigma_{x})]^{-1}$ that acts to approximately whiten the input data $x_{q}$ relative to the context statistics $\hat{\Sigma}_{x}$. This operator effectively inverts the data covariance modulated by finite-context effects ($N$), allowing the model to isolate the underlying linear relationship. Broadly, our work demonstrates how analytical techniques from deep linear network theory [10] can be extended to analyze context-dependent computations in transformers, revealing that timescale separation can arise directly from the interaction of attention parameters with data statistics, not just from network depth. Furthermore, the analysis revealed aspects of low-dimensional learning. First, parameter specialization occurs, where certain parameter blocks ($p_{1},q_{2}$) remain "null" while others ($p_{2},q_{1}$) actively drive learning. Second, a conservation law constrains the dynamics, reducing the effective dimensionality of the parameter trajectory and highlighting symmetries in the learning landscape. Crucially, the qualitative features predicted by our solvable model—timescale separation, sequential mode learning, subspace stabilization—appear empirically relevant in non-linear attention-only models and grokking scenarios. The alignment between theoretical predictions and empirical observations using measures like effective rank and subspace distance suggests that the core mechanisms identified in the linear model offer valuable insights into the learning processes of more complex transformers.
Finally, this work suggests a principled approach to interpretability and evaluation. The theoretically grounded measures developed here—loss/norm curvature, marginalized effective rank, subspace distance—can empirically identify critical time points during training associated with significant learning events or potential shifts in model capabilities. Monitoring these metrics could provide a more informed strategy for timing behavioral evaluations, potentially capturing the emergence of new functionalities more effectively than relying solely on overall loss or accuracy measures. This offers a promising direction for bridging theoretical understanding with practical monitoring and analysis of large model training. Several avenues remain for future investigation. Incorporating non-linear attention mechanisms is essential to understand their impact on dynamics and potential deviations from the linear theory. Non-Gaussian data model can give rise to richer dynamics due to non-trivial higher moments and might lead to better predictions on dynamics in real-world models. Analyzing the effects of network depth and multiple attention heads could reveal how scaling influences ICL capabilities and learning complexity. Extending the analysis beyond linear regression to tasks involving non-linear functions or discrete structures is also a critical next step.
@@ A.7 Key Assumptions
  • Prediction focused on the $(\bm{N+1})$-th output token. We focus our analysis on the prediction derived from the $(N+1)$-th output token, corresponding to the query input $x_{q}$.
@@ A.11.1 SGD Update for $p_{2}$
$\frac{\partial h(p_{2})}{\partial p_{2}}=\frac{\partial}{\partial p_{2}}\left[p_{2}W^{\mu}\hat{\Sigma}_{x}q_{1}x_{q}-W^{\mu}x_{q}\right]=\frac{\partial}{\partial p_{2}}\left[p_{2}(W^{\mu}\hat{\Sigma}_{x}q_{1}x_{q})-W^{\mu}x_{q}\right]=x_{q}^{\top}q_{1}^{\top}\hat{\Sigma}_{x}(W^{\mu})^{\top},$ (A57)
@@ A.11.2 Change of Variables and Assumptions for Tractable Dynamics
$=-\lambda\sum_{\mu=1}^{P}\left(\bar{p}_{2}\Lambda^{\mu}\hat{S}^{2}\bar{q}_{1}^{2}\bar{x}_{q}\bar{x}_{q}^{\top}\Lambda^{\mu}-\Lambda^{\mu 2}\bar{q}_{1}\bar{x}_{q}\bar{x}_{q}^{\top}\hat{S}\right),$ (A71)
@@
$\mathbb{E}_{\Lambda}[\Delta\bar{p}_{2}]=-\lambda P\left(\bar{p}_{2}\hat{S}^{2}\bar{q}_{1}^{2}\bar{x}_{q}\bar{x}_{q}^{\top}\mathbb{E}_{\Lambda}[\Lambda^{2}]-\mathbb{E}_{\Lambda}[\Lambda^{2}]\bar{q}_{1}\bar{x}_{q}\bar{x}_{q}^{\top}\hat{S}\right).$ (A73)
@@
$\mathbb{E}_{\Lambda}[\Delta\bar{p}_{2}]=-\lambda P\left(\bar{p}_{2}\hat{S}^{2}\bar{q}_{1}^{2}\bar{x}_{q}\bar{x}_{q}^{\top}-\bar{q}_{1}\bar{x}_{q}\bar{x}_{q}^{\top}\hat{S}\right).$ (A74)
@@
$\mathbb{E}_{\Lambda,x_{q}}[\Delta\bar{p}_{2}]=-\lambda P\left(\bar{p}_{2}\hat{S}^{2}\bar{q}_{1}^{2}S-\bar{q}_{1}S\hat{S}\right)=-\lambda PS\left(\bar{p}_{2}\hat{S}^{2}\bar{q}_{1}^{2}-\bar{q}_{1}\hat{S}\right).$ (A75)
@@
$\mathbb{E}[\Delta\bar{p}_{2}]$ $=-\lambda PS\left(\bar{p}_{2}\bar{q}_{1}^{2}\mathbb{E}[\hat{S}^{2}]-\bar{q}_{1}\mathbb{E}[\hat{S}]\right)$ (A76)
@@ A.12.2 Diagonal Element Dynamics and Solution
[…]
@@
under the assumption $p_{\alpha}=q_{\alpha}$, so $p_{\alpha}^{2}=q_{\alpha}^{2}=a_{\alpha}$. The fixed point for $a_{\alpha}$ is obtained by setting $\frac{da_{\alpha}}{dt}=0$, which gives $02a_{\alpha}\left(1-a_{\alpha}s_{\alpha}^{\infty}\right)=$ 0. Non-trivial fixed point is $a_{\alpha}^{\infty}=\frac{1}{s_{\alpha}^{\infty}}=\frac{1}{\frac{N+1}{N}s_{\alpha}+\frac{\operatorname{Tr}(S)}{N}}$.
@@
[…]
@@
[…]

### v2

@@ 4 Applications to Analysis of Non-linear Transformer Models
Probing Timescale Separation via Curvature. The solved loss dynamics (Eq. (14)) predict non-uniform learning rates across different spectral modes $\alpha$, leading to characteristic plateaus followed by cliffs. Detecting these transitions in the raw training loss of non-linear models is challenging due to the stochastic noise inherent in SGD, which obscures subtle inflection points. To mitigate this, we analyze the curvature of the loss autocorrelation function, defined as $A(\tau)=\mathbb{E}[L(t)L(t+\tau)]$. This metric acts as a smoothing filter, preserving macroscopic trend changes while suppressing high-frequency batch noise. Peaks in the second derivative of $A(\tau)$ indicate rapid changes in the rate of loss decay, corresponding to the ”cliffs” where a new eigenmode is rapidly learned (Figure 4A). Complementarily, we analyze the curvature of the Frobenius norm of the active parameter matrices ($\|p_{2}(t)\|_{F}$ or $\|q_{1}(t)\|_{F}$). Since the parameters $p_{\alpha},q_{\alpha}$ evolve according to Eq. (12) with mode-dependent rates, their overall norm should also exhibit non-uniform changes. Peaks in the parameter norm curvature align well with those in the loss autocorrelation (Figure 4B), reinforcing that macroscopic observables reflect the underlying timescale separation predicted by the theory.
@@
Tracking Sequential Learning via Effective Rank. The exact solution $a_{\alpha}(t)$ (Eq. (13)) describes the sequential ”activation” or learning of different spectral modes. To track this process empirically in the weight matrices of non-linear models, we utilize the effective rank (roy2007effective), which measures the dimensionality or concentration of singular values ($s_{i}$) of a matrix $M$:
@@
A lower effective rank implies energy concentration in fewer singular modes. Motivated by the theoretical prediction that dominant modes (large $s_{\alpha}$) learn first (small $\tau_{\alpha}$), we expect the effective rank to initially decrease as the model focuses on these modes, and subsequently increase as subdominant modes are learned. To better visualize this sequential process, we introduce ”marginalized effective rank”. This involves computing the effective rank repeatedly, each time excluding the singular value component corresponding to the mode assumed to have converged earliest. Dips in these marginalized curves should indicate the time points where specific spectral components undergo significant learning (Figure 4C). As predicted by the sequential learning in Eq. (13), the effective rank initially dips, then recovers, and the marginalized curves reveal staggered learning across modes.
@@
We applied these analyses to attention-only non-linear transformers (1-4 layers), similar to those exhibiting ICL capabilities (olsson2022context). Figure 6 displays the marginalized effective rank for Output-Value (OV) and Query-Key (QK) matrices. Consistent with the timescale separation hypothesis, models capable of ICL (2+ layers) show characteristic dips in the effective rank curves, particularly for QK matrices, around the time ICL performance emerges (highlighted region). The full set of marginalized curves reveals multiple minima occurring sequentially, supporting the idea derived from our linear model that learning proceeds in stages along different spectral dimensions. In contrast, the 1-layer model, which fails at this ICL task, shows a monotonic increase in OV effective rank, suggesting undifferentiated learning or noise fitting rather than structured, sequential learning. These empirical results suggest that the sequential, mode-dependent learning predicted by our solvable model qualitatively extends to these non-linear settings and correlates with the emergence of functional capabilities like ICL.
Stability of Parameter Subspace. Our theoretical analysis separates the dynamics of parameter magnitudes (related to $a_{\alpha}=p_{\alpha}q_{\alpha}$) from the underlying parameter structure. The decoupled dynamics (Eq. (12)) allow for the possibility that the directions spanned by the singular vectors of weight matrices stabilize faster than the singular values reach their fixed points. To test this, we introduce a ”Subspace Distance” metric, measuring the alignment between the subspace spanned by the singular vectors of a weight matrix $M(t)$ at time $t$ and its final state $M(\infty)$:
$\textrm{Sub. Dist.}\left[M(t)\right]=\min_{A\in\mathbb{R}^{D\times D}}\left\|AM(t)-M(\infty)\right\|.$ (16)
@@
This hypothesis is further tested by examining models exhibiting “grokking” in modular arithmetic tasks (nanda2023grokking), where generalization is significantly delayed relative to training performance saturation. Figure 3B shows the characteristic delayed drop in test error. Analyzing the OV matrix dynamics, we observe that subspace stabilization, measured by the Subspace Distance (Figure 7C), is markedly delayed. It occurs significantly after the training loss has minimized and coincides closely with the sharp improvement in generalization (test error drop). The effective rank (Figure 7B) also shows dynamics linked to this delayed phase. This finding strongly suggests that grokking, in this context, is linked to a delayed convergence of the parameter subspace, providing a potential mechanistic explanation. Once the correct subspace is identified (albeit late), learning of magnitudes proceeds rapidly, leading to generalization.
@@ 5 Discussion
The learned computation at the fixed point (Eq. (11)) effectively implements the target linear map $W^{\mu}$ by applying an operator $\hat{\Sigma}_{x}[s^{\infty}(\Sigma_{x})]^{-1}$ that acts to approximately whiten the input data $x_{q}$ relative to the context statistics $\hat{\Sigma}_{x}$. This operator effectively inverts the data covariance modulated by finite-context effects ($N$), allowing the model to isolate the underlying linear relationship. Broadly, our work demonstrates how analytical techniques from deep linear network theory (saxe2014exact) can be extended to analyze context-dependent computations in transformers, revealing that timescale separation can arise directly from the interaction of attention parameters with data statistics, not just from network depth. Furthermore, the analysis revealed aspects of low-dimensional learning. First, parameter specialization occurs, where certain parameter blocks ($p_{1},q_{2}$) remain ”null” while others ($p_{2},q_{1}$) actively drive learning. Second, a conservation law constrains the dynamics, reducing the effective dimensionality of the parameter trajectory and highlighting symmetries in the learning landscape. Crucially, the qualitative features predicted by our solvable model—timescale separation, sequential mode learning, subspace stabilization—appear empirically relevant in non-linear attention-only models and grokking scenarios. The alignment between theoretical predictions and empirical observations using measures like effective rank and subspace distance suggests that the core mechanisms identified in the linear model offer valuable insights into the learning processes of more complex transformers.
Finally, this work suggests a principled approach to interpretability and evaluation. The theoretically grounded measures developed here—loss/norm curvature, marginalized effective rank, subspace distance—can empirically identify critical time points during training associated with significant learning events or potential shifts in model capabilities. Monitoring these metrics could provide a more informed strategy for timing behavioral evaluations, potentially capturing the emergence of new functionalities more effectively than relying solely on overall loss or accuracy measures. This offers a promising direction for bridging theoretical understanding with practical monitoring and analysis of large model training. Several avenues remain for future investigation. Incorporating non-linear attention mechanisms is essential to understand their impact on dynamics and potential deviations from the linear theory. Analyzing the effects of network depth and multiple attention heads could reveal how scaling influences ICL capabilities and learning complexity. Extending the analysis beyond linear regression to tasks involving non-linear functions or discrete structures is also a critical next step.
@@ A.7 Key Assumptions
  • Prediction focused on the $(\boldsymbol{N+1})$-th output token. We focus our analysis on the prediction derived from the $(N+1)$-th output token, corresponding to the query input $x_{q}$.
@@ A.11.1 SGD Update for $p_{2}$
$\frac{\partial h(p_{2})}{\partial p_{2}}=\frac{\partial}{\partial p_{2}}\left[p_{2}W^{\mu}\hat{\Sigma}_{x}q_{1}x_{q}-W^{\mu}x_{q}\right]=\frac{\partial}{\partial p_{2}}\left[p_{2}(W^{\mu}\hat{\Sigma}_{x}q_{1}x_{q})-W^{\mu}x_{q}\right]=(W^{\mu}\hat{\Sigma}_{x}q_{1}x_{q})^{\top}=x_{q}^{\top}q_{1}^{\top}\hat{\Sigma}_{x}^{\top}(W^{\mu})^{\top}=x_{q}^{\top}q_{1}^{\top}\hat{\Sigma}_{x}(W^{\mu})^{\top},$ (A57)
@@ A.11.2 Change of Variables and Assumptions for Tractable Dynamics
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ A.12.2 Diagonal Element Dynamics and Solution
Observe that $\tau_{\alpha}p_{\alpha}\frac{dp_{\alpha}}{dt}=p_{\alpha}q_{\alpha}\left(1-p_{\alpha}q_{\alpha}s_{\alpha}^{\infty}\right)$ and $\tau_{\alpha}q_{\alpha}\frac{dq_{\alpha}}{dt}=p_{\alpha}q_{\alpha}\left(1-p_{\alpha}q_{\alpha}s_{\alpha}^{\infty}\right)$. Thus, $\tau_{\alpha}p_{\alpha}\frac{dp_{\alpha}}{dt}=\tau_{\alpha}q_{\alpha}\frac{dq_{\alpha}}{dt}$, which implies $p_{\alpha}\frac{dp_{\alpha}}{dt}=q_{\alpha}\frac{dq_{\alpha}}{dt}$, or $p_{\alpha}dp_{\alpha}=q_{\alpha}dq_{\alpha}$. Integrating both sides, we get $\int p_{\alpha}dp_{\alpha}=\int q_{\alpha}dq_{\alpha}$, leading to $\frac{1}{2}p_{\alpha}^{2}=\frac{1}{2}q_{\alpha}^{2}+C^{\prime}_{\alpha}$, or $p_{\alpha}^{2}-q_{\alpha}^{2}=C_{\alpha}$, where $C_{\alpha}=2C^{\prime}_{\alpha}$ is a constant of integration. This is the element-wise conserved quantity, consistent with the matrix conserved quantity $\mathcal{C}$. If initialized with $p_{\alpha}(0)=q_{\alpha}(0)$, then $C_{\alpha}=0$, and $p_{\alpha}^{2}(t)=q_{\alpha}^{2}(t)$ for all $t$. Assuming $p_{\alpha}(t)$ and $q_{\alpha}(t)$ maintain the same sign (if initialized non-negative, remain non-negative), we can consider $p_{\alpha}(t)=q_{\alpha}(t)$.
@@
under the assumption $p_{\alpha}=q_{\alpha}$, so $p_{\alpha}^{2}=q_{\alpha}^{2}=a_{\alpha}$. The fixed point for $a_{\alpha}$ is obtained by setting $\frac{da_{\alpha}}{dt}=0$, which gives $2a_{\alpha}\left(1-a_{\alpha}s_{\alpha}^{\infty}\right)=0$. Non-trivial fixed point is $a_{\alpha}^{\infty}=\frac{1}{s_{\alpha}^{\infty}}=\frac{1}{\frac{N+1}{N}s_{\alpha}+\frac{\operatorname{Tr}(S)}{N}}$.
@@
$\int\frac{da_{\alpha}}{a_{\alpha}\left(1-a_{\alpha}s_{\alpha}^{\infty}\right)}=\int\frac{2}{\tau_{\alpha}}dt.$
@@
$\int\left(\frac{1}{a_{\alpha}}+\frac{s_{\alpha}^{\infty}}{1-a_{\alpha}s_{\alpha}^{\infty}}\right)da_{\alpha}$ $=\frac{2t}{\tau_{\alpha}}+C^{\prime\prime}_{\alpha}$